A particle has acceleration (a) while moving with uniform speed v in a circle of radius r. Start with a rnvm. Derive the relationship between a, r and v using dimensional analysis. 


1
Expert's answer
2021-03-02T18:07:26-0500

Explanations & Calculations


  • To deal with these questions, first, try to write the units of all the given quantities (on both sides of the equal mark) in their SI units.
  • Then gather like units together & compare their indexes.

"\qquad\qquad
\begin{aligned}
\small a&\propto r^n\cdot v^m\\
\small [a]&= \small LT^{-2}\\
\small [r]&= \small L\\
\small [v]&= \small LT^{-1}\\
\\
\small LT^{-2}&\propto \small L^n\cdot (LT^{-1})^m\\
&\propto \small L^{(n+m)}\cdot T^{-m}\\
\\
\small L:\,\,\,\qquad1&=n+m\cdots(1)\\
\small T:\qquad-2&\propto \small -m\\
\\
\small m&= \small 2\\
\small n&= \small 1-2=-1
\end{aligned}"

  • Then what we are able to write is

"\qquad\qquad
\begin{aligned}
\small a&\propto \small k\cdot\frac{v^2}{r}\qquad or\qquad a= k\cdot\frac{v^2}{r}
\end{aligned}"

  • The dimensional analysis only allows us to estimate the validity of an equation for a relationship, not fully derive the real relationship.
  • For that experimental approach is needed.